ND-GAr: HPgTPC+ECAL Weekly Meeting

Disappearance Room (WH13SE) and Zoom

Disappearance Room (WH13SE) and Zoom

    • 10:00 AM 10:20 AM
      Roundtable Updates 20m
      Speakers: Alan Bross (Fermilab), Alfons Weber (JG-Uni Mainz & Fermilab), Diego Gonzalez Diaz, Naseem Khan, Tanaz Mohayai (Fermilab), vivek jain (SUNY Albany)

      Collaboration meeting:  please let us know if you want to give a talk.  Asa thinks that they won't be able to give a talk.

      GORG:   Tanaz currently taking data at 1 atm. 

      TOAD:  Naseem:  Replacement for third agregator has  arrived.  Trying to improve the grounding.  Hope to be able to fill soon to 1 bar.  Ran vacuum all weekend.  Took beam data yesterday with all four wire chambers and the magnet.  All four chambers work more or less.  Now beam reaches 3 and 4 with the magnet on.  Still working on trying to get time of flight working with DAQ.    Anezka:  Naseem can now decode HFA files and has something that can hopefully convert to root.  Once she had files can try them out in GArsoft and run in event display.  Alan asked about trigger.  Will feed in the $30 start of spill and also will try to feed in beamline trigger from ToF counters.

      GOAT:  Alan showed some plots of GOAT data at 10 bar.  But all data with various CH4 content on one plot.    In PEP4 gain was only 1000 with 3700 V, but it was an 80/20 mix to reduce diffusion.    Would be nice in TOAD to maybe try to dilute gas to get more gain.    Default for TOAD is likely 96%/4%.  Alan suggests that 98%/2% might be a bit better and at the plateau in drift voltage. 

      ECAL:   No updates. But Vivek hopes to make more progress soon.  Wants to understand geometry code and why it is giving wrong information.